If 2.33 g of CuNO3 is dissolved in water to make a 0.250 M solution, what is the volume of the solution in millilitres


1
Expert's answer
2021-12-21T11:11:03-0500

The molar mass of CuNO3 is 125.54 g/mol

"V_{solution}=2.33\ g(CuNO_3)\times\frac{1\ mol}{125.54\ g(CuNO_3)}\times\frac{1\ L}{0.250\ mol}\times\frac{1000\ mL}{1\ L}=74.2\ mL"


Answer: 74.2 mL